OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 9 when x = 5. What is the constant of variation in decimal form?

TheSmartOne (thesmartone):

\(\sf\Large k = \frac{y}{X}\)

OpenStudy (anonymous):

...

TheSmartOne (thesmartone):

So plug in the values of x and y into it...

TheSmartOne (thesmartone):

and k is the direct variation which would be your answer.

OpenStudy (anonymous):

ok, so it would look like this? \[k=\frac{ 5 }{ 9 }\]?

TheSmartOne (thesmartone):

\(\sf\Huge\frac{y}{x}\) not \(\sf\Large\frac{x}{y}\)

OpenStudy (anonymous):

my bad \[k=\frac{ 9 }{ 5 }\]

OpenStudy (anonymous):

BOI YOU LEFT ME

TheSmartOne (thesmartone):

And just make that into a decimal which is 1.8

OpenStudy (anonymous):

ok, ^_^ thanks

TheSmartOne (thesmartone):

I'm sorry that I have so many tags and notifications :P Currently 70 atm :P

OpenStudy (anonymous):

Ohhh, busy guy

OpenStudy (anonymous):

or ggggirl

TheSmartOne (thesmartone):

guy :p